A hockey puck B rests in a smooth ice surface and is struck by a second puck A,



which was originally traveling at 40.0 m/s and is deflected 30.0° from its original



direction. The pucks have the same mass. (a) Compute the speed of each puck after



the collision. (b) What fraction of the original kinetic energy of puck A dissipates



during collision?

1
Expert's answer
2022-01-31T11:25:10-0500

Conservation of momentum by axes:


"0=mu_A\sin\theta_A+mu_B\sin\theta_B,\\
mv_A=mu_A\cos\theta_A+mu_B\cos\theta_B.\\\space\\

0=u_A\sin\theta_A+u_B\sin\theta_B,\\
v_A=u_A\cos\theta_A+u_B\cos\theta_B.\\\space\\

u_A =29.3 \text{ m/s},\\
u_B =20.1 \text{ m/s}."

The fraction of energy:


"ϵ_E =\dfrac{v_A^2 −u_A^2}{v_A^2} =0.46."